OpenStudy (sissyedgar):

ASAP HELP PLEASE!!! MEDALS! Evaluate the expression when t = –6.4 and u = –2.8. t – 4u A. –17.6 B. 4.8 C. 17.6 D. 29.12

OpenStudy (sissyedgar):

@cheerleader

OpenStudy (anonymous):

B

OpenStudy (cheerleader):

its d babe

OpenStudy (anonymous):

make t= to -6.4 and u=-2.8 -6.4-4(-2.8)=?

OpenStudy (anonymous):

umm no its nnot

OpenStudy (anonymous):

its b

OpenStudy (sissyedgar):

thx

OpenStudy (anonymous):

i did it in the calc to double cheak its 4.8

OpenStudy (anonymous):

stop just saying the answers guys, you have to explain how to get the answer, or she will not learn anything.

OpenStudy (sissyedgar):

i am in a hurrrry, i will learn later, so d or b

OpenStudy (anonymous):

b

OpenStudy (anonymous):

its really easy to learn though i could teach you now

OpenStudy (sissyedgar):

its D

OpenStudy (anonymous):

no b

OpenStudy (sissyedgar):

i think

OpenStudy (sissyedgar):

i am checkin now. i passed doe

OpenStudy (anonymous):

its B...

OpenStudy (anonymous):

-6.4-4(-2.8)=4.8

OpenStudy (sissyedgar):

SORRY!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

lol no my mistake its d

OpenStudy (sissyedgar):

no. its b

OpenStudy (sissyedgar):

cordin 2 the assighnment

OpenStudy (anonymous):

lol im so stupid

OpenStudy (sissyedgar):

u think u can help me wiff 5 more?